OpenStudy (anonymous):

A plane bisects a 90 degree dihedral angle. From a point on this plane 16 in. from the common edge, perpendicular lines are constructed to the respective faces of the dihedral angle. Find the length of each perpendicular

OpenStudy (trantom):



OpenStudy (trantom):

so that's what we're looking at and we need to find AB and AD (but really just one of them because they are congruent)

OpenStudy (trantom):

well if we look at just the triangle ABC, we know that the two legs are congruent because those two original planes were bisected so you use the Pythagorean therom 16^2=2*a^2 256 = 2 * a^2 divide by 2 128 = a^2 sqrt(128) = a so there you go the answer is the square root of 128